### v1

# Mutual Information Free Topological Generalization Bounds via Stability

Abstract
Providing generalization guarantees for stochastic optimization algorithms is a major challenge in modern learning theory. Recently, several studies highlighted the impact of the geometry of training trajectories on the generalization error, both theoretically and empirically. Among these works, a series of topological generalization bounds have been proposed, relating the generalization error to notions of topological complexity that stem from topological data analysis (TDA). Despite their empirical success, these bounds rely on intricate information-theoretic (IT) terms that can be bounded in specific cases but remain intractable for practical algorithms (such as ADAM), potentially reducing the relevance of the derived bounds. In this paper, we seek to formulate comprehensive and interpretable topological generalization bounds free of intractable mutual information terms. To this end, we introduce a novel learning theoretic framework that departs from the existing strategies via proof techniques rooted in algorithmic stability. By extending an existing notion of hypothesis set stability, to trajectory stability, we prove that the generalization error of trajectory-stable algorithms can be upper bounded in terms of (i) TDA quantities describing the complexity of the trajectory of the optimizer in the parameter space, and (ii) the trajectory stability parameter of the algorithm. Through a series of experimental evaluations, we demonstrate that the TDA terms in the bound are of great importance, especially as the number of training samples grows. This ultimately forms an explanation of the empirical success of the topological generalization bounds.
1 Introduction
The generalization capabilities of deep learning models play a crucial role in our understanding of modern machine learning. Explaining generalization and, thus, the success of current neural networks is still an open problem and an active research area [ZBH+17, ZBH+21]. Following classical statistical learning theory [SSBD14], supervised learning tasks can be framed as a population risk minimization problem, defined as:
$\min_{w\in\mathbb{R}^{d}}\left\{\mathcal{R}(w):=\mathbb{E}_{z\sim\mu_{z}}[\ell(w,z)]\right\},$ (1)
where $z\in\mathcal{Z}$ denotes the data, distributed according to a probability distribution $\mu_{z}$ on the data space $\mathcal{Z}$ and $\ell:\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}$, the composed loss function. In practice, we have $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ denotes a feature space and $\mathcal{Y}$ is a label space. Since the data distribution $\mu_{z}$ is unknown, it is in general not possible to directly solve (1); instead, we typically solve the empirical risk minimization problem (2): for $n$ given data points $S:=(z_{1},\dots,z_{n})\sim\mu_{z}^{\otimes n}:=\mu_{z}\otimes\dots\otimes\mu_{z}$ independently sampled from the data distribution $\mu$, we consider:
$\min_{w\in\mathbb{R}^{d}}\left\{\widehat{\mathcal{R}}_{S}(w):=\frac{1}{n}\sum_{i=1}^{n}\ell(w,z_{i})\right\}$ (2)
In most applications, the problem (2) is solved by the means of stochastic optimization algorithms, such as stochastic gradient descent (SGD) or ADAM [KB14]. These algorithms often take an iterative form and we denote $w_{k}$ the parameter that is learned after $k$ iterations.
Recent studies suggest that the training trajectory, i.e., the set of iterates $w_{k}$ obtained by the algorithm, may be paramount in explaining the generalization ability of modern machine learning models [XTYY24, FZY+24, ADS+24]. Indeed, since the empirical risk (2) can contain many (typically infinitely many) local minima with different generalization properties, the trajectories of the optimization algorithm near a local minimum offer a concise way to assess the quality of the local minimum: they inherently reflect the influence of the optimization algorithm, the choice of hyperparameters, and the dataset—factors that are all essential for deriving meaningful generalization bounds.
More precisely, the trajectory of the optimizer is essentially a random subset of $\mathbb{R}^{d}$, that is defined as follows: $\mathcal{W}_{S,U}:=\{w_{t}:t_{0}\leq t\leq T\}\subset\mathbb{R}^{d}$, where $t_{0}$ is chosen such that $w_{t_{0}}$ is near a local minimum that is found by the optimizer, say $w_{\star}$, $U$ represents the algorithmic randomness (e.g., batch indices), which is independent of the other random variables. In essence, $\mathcal{W}_{S,U}$ is the part of the basin corresponding to $w_{\star}$ that is accessible to the optimization algorithm. To measure the quality of the local minimum $w_{\star}$ by using the structure of the trajectory $\mathcal{W}_{S,U}$, the main quantity of interest has been the worst-case generalization error over $\mathcal{W}_{S,U}$ , defined as follows:
$G_{S}(\mathcal{W}_{S,U}):=\sup_{w\in\mathcal{W}_{S,U}}\left(\mathcal{R}(w)-\hat{\mathcal{R}}_{S}(w)\right).$ (3)
Upper-bounding $G_{S}(\mathcal{W}_{S,U})$ requires uniform convergence results on the random, data-dependent hypothesis set $\mathcal{W}_{S,U}$, which is a notoriously challenging problem [FGK+19, SvEH+23, DVDS24].
In this context, recent studies [SSDE20, DDS23, ALRS23, HSKM22, DVDS24, BLGS21, ADS+24] have established upper-bounds on (3) and drawn rigorous links between the topological structure of $\mathcal{W}_{S,U}$ and $G_{S}(\mathcal{W}_{S,U})$. These bounds are of the following form [^1] We use the notation $\lesssim$ in informal statements where absolute constants have been omitted.:
$G_{S}(\mathcal{W}_{S,U})\lesssim\sqrt{\frac{\mathbf{C}(\mathcal{W}_{S,U})+\mathrm{IT}+\log(1/\zeta)}{n}},$ (4)
with probability at least $1-\zeta$. The bound in (4) has two main ingredients: (i) The first term $\mathbf{C}(\mathcal{W}_{S})$, which we call the topological complexity in accordance with [ADS+24], is a geometric quantity of the trajectory, measuring some notion of complexity of $\mathcal{W}_{S}$. Several topological complexities have been proposed, inspired by fractal geometry [SSDE20, DDS23, DV23] and topological data analysis (TDA) [BLGS21, ALRS23, ADS+24]. (ii) The second term, $\mathrm{IT}$, is an information-theoretic (IT) term, typically the total mutual information between the dataset $S$ and the trajectory $\mathcal{W}_{S}$; measuring the statistical dependence between $S$ and $\mathcal{W}_{S,U}$ [DVDS24].
The topological complexity term $\mathbf{C}(\mathcal{W}_{S,U})$ appearing in these bounds are both computable and practically informative. In contrast, the $\mathrm{IT}$ term remains opaque and computationally intractable; its behavior with respect to the sample size $n$ is not even well understood in the general case [DVDS24]. As a result, prior studies have adopted a heuristic approach by entirely discarding the IT term in their empirical evaluations and concentrating solely on $\mathbf{C}(\mathcal{W}_{S,U})$. These studies reported striking empirical results, demonstrating a strong correlation between the generalization error and the proposed topological complexity measures (an observation we will also report in our experiments) [^2] While [TGRW+24] recently raised some concerns about the practical validity of certain fractal dimension–based bounds, [ADS+24] showed that specific topological descriptors—such as positive magnitude and $\alpha$-weighted lifetime sums, which we will detail in the next section, consistently yield strong empirical performance across a wide range of modern deep learning scenarios.. However, the exclusion of the IT term renders such evaluations theoretically incomplete, and the underlying reason for the empirical success of the topological complexity measures remains unclear.
To address this issue, an initial step was taken in [SvEH+23], where the authors introduced the notion of ‘algorithm-dependent Rademacher complexity’ (ARC) and derived generalization bounds with a topological flavor, notably avoiding the $\mathrm{IT}$ term. However, ARC is based on combining parameters in a combinatorial fashion (similar to the approach that has been used for conditional mutual information-based bounds [SZ20]). As a result, the associated topological quantities were not tied to the data-dependent optimization trajectory $\mathcal{W}_{S,U}$, but rather to an abstract, theoretically constructed hypothesis set. Consequently, while the bounds are insightful from a theoretical standpoint, they fall short of explaining why $\mathbf{C}(\mathcal{W}_{S,U})$ proves effective in practice.
In this study, our goal is to bridge the gap between topological generalization bounds and the aforementioned empirical observations, by deriving generalization bounds that are directly expressed in terms of the $\mathbf{C}(\mathcal{W}_{S,U})$ while entirely avoiding the information-theoretic terms.
Contributions
Our main contributions are summarized below.
  • Inspired by [FGK+19], we define the notion of trajectory stability, that is adapted to stochastic learning algorithms, i.e., producing random trajectories. Based on this stability condition, we prove expected topological generalization bounds of the form:
    $\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\left(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\right)\right]\lesssim\beta_{n}^{1/3}\left(1+\mathbb{E}\left[\sqrt{\mathbf{C}(\mathcal{W}_{S,U})}\right]\right),$
    where $\beta_{n}$ is the stability parameter and $\mathbf{C}(\mathcal{W}_{S,U})$ is a topological complexity identical to some of the TDA-inspired terms used in [BLGS21, ADS+24]. In particular, we recover the so-called $\alpha$-weighted lifetime sums and positive magnitude based on the Euclidean distance (see Section 2).
  • We further investigate our new trajectory stability condition and show that a variant of the classical notion of algorithmic stability [HRS16] implies trajectory stability. Hence, if the optimization algorithm is algorithmically stable in a more classical sense, existing algorithmic stability results imply that, in various cases, the trajectory stability parameter $\beta_{n}$ will be of order $\mathcal{O}(1/n)$, rendering our bounds $\mathcal{O}(n^{-1/3}\sqrt{\mathbf{C}(\mathcal{W}_{S})})$.
  • We extend recent empirical studies on topological generalization bounds and explain the interplay between stability and topology, emphasizing the role of sample size as a key factor in the analysis of topological generalization bounds.
Our results demonstrate that, if the optimizer is trajectory-stable, the information-theoretic terms in the bound Equation 4 can be safely omitted, at the expense of a reduced convergence rate [^3] A similar convergence rate was also reported in [DDS23] in an attempt to reduce the impact of the IT terms., i.e., from $n^{-1/2}$ to $n^{-1/3}$. We believe that our findings reduce the gap between theory and practice, paving the way for even more practical topological generalization bounds with new empirical insights for large-scale datasets. The implementation will be made publicly available at https://github.com/bastianlb/MI-free-topo-bounds. All the proofs are provided in the appendix.
2 Preliminaries Technical Background
We now introduce technical background regarding the topological complexities used in this work (Section 2.1) and algorithmic stability (Section 2.2).
2.1 Topological Complexity Measures
$\alpha$-Weighted lifetime sum
The $\alpha$-weighted lifetime sum is one of the building blocks of the persistent homology dimension [Sch20, BLGS21]. Lifetime intervals encode how long topological features (like connected components) persist across scales [BGND+19, HKNU17, HKND19]. Informally, it tracks a finite set’s “connected components” at different scales. For simplicity and readability, we provide an equivalent formulation based on minimum spanning trees [Sch20, KLS06].
Let $X$ be a finite set of points endowed with a metric $\rho$. A tree over X is a connected acyclic undirected graph whose vertices are the points in X. Given an edge $e$ linking the points a and b, we define its cost as $|e|:=\rho(a,b)$ The $\alpha$-weighted lifetime sums are defined as:
$\forall\alpha\geq 0,\quad\textbf{E}_{\rho}^{\alpha}(X):=\sum_{e\in\mathrm{MST}(X)}|e|^{\alpha},$
where $\mathrm{MST}(X)$ denotes a minimum spanning tree over the graph defined by $X$.
Positive magnitude
The concept of metric magnitude was introduced in [Lei13], it is critical in geometric measure theory and integral geometry [Mec13]. In [ADS+24] a variant called positive magnitude was introduced and rigorously linked to generalization error. Informally, (positive) magnitude can be interpreted as the effective number of distinct points in a space at a given scale. We now give the definition of positive magnitude. For a “scale” parameter $s>0$ and a finite metric space $(X,\rho)$, we define a weighting of the scaled space $(s\cdot X,\rho)$ as a vector $\gamma_{s}:X\to\mathbb{R}$ such that for any $a\in X$, we have $\sum_{b\in X}e^{-s\rho(a,b)}\gamma_{s}(b)=1$. The existence of such weighting is ensured in the finite Euclidean setting [Mec15]. The positive magnitude of $(s\cdot X,\rho)$ is then defined as:
$\mathbf{PMag}_{\rho}(s\cdot X):=\sum_{b\in X}\gamma_{s}(b)_{+},$ (5)
where $\gamma_{+}:=\max\{0,\gamma\}$. The parameter $s$ can be seen as a scale through which we look at $(X,\rho)$.
2.2 Algorithmic stability
Algorithmic stability is a fundamental concept in learning theory [BE02, BKZ20]. It has proven successful in providing generalization guarantees for widely used stochastic optimization algorithms [BFGT20, FV19, ZGRS23, HRS16]. In this paper, we consider a variant of algorithmic stability, called uniform argument stability, introduced by [BFGT20].
In all the following, we say that two datasets $S,S^{\prime}\in\mathcal{Z}^{n}$ are neighboring datasets if they differ by one element, which we denote $S\simeq S^{\prime}$.
Definition 2.1 (Uniform argument stability [BFGT20]).
Let $\mathcal{A}(S,U)\in\mathbb{R}^{d}$ be a randomized algorithm where $S\in\mathcal{Z}^{n}$ is a dataset and $U$ represent the algorithm randomness. We say that $\mathcal{A}$ satisfies uniform parameter stability with coefficient $\beta>0$ if for any neighboring datasets $S\simeq S^{\prime}$, we have:
$\mathbb{E}_{U}\left[\left\|\mathcal{A}(S,U)-\mathcal{A}(S^{\prime},U)\right\|\right]\leq\beta.$
If $\ell(w,z)$ is $L$-Lipschitz continuous in $w$, this implies the classical algorithmic stability [HRS16]:
$\sup_{S\simeq S^{\prime}}\sup_{z\in\mathcal{Z}}\mathbb{E}_{U}\left[\ell(\mathcal{A}(S,U),z)-\ell(\mathcal{A}(S^{\prime},U),z)\right]\leq L\beta.$ (6)
Unfortunately, in the context of worst case generalization bounds (i.e., upper bounds on $G_{S}(\mathcal{W}_{S})$), this assumption is not of much use as one needs to specify the stability of the entire set $\mathcal{W}_{S}$ and not just one parameter inside it. To address this issue, a first step was taken by [FGK+19], who introduced the following notion of hypothesis set stability.
Definition 2.2 ([FGK+19]).
A family $(\mathcal{W}_{S})_{S\in\mathcal{Z}^{n}}$ of data-dependent hypothesis sets is said to be $\beta$-uniformly stable if for any $S\simeq S^{\prime}$ in $\mathcal{Z}^{n}$ and any $w\in\mathcal{W}_{S}$, if there exists $w^{\prime}\in\mathcal{W}_{S^{\prime}}$ such that:
$\sup_{z\in\mathcal{Z}}|\ell(w,z)-\ell(w^{\prime},z)|\leq\beta.$
Note that the classical algorithmic stability (as described in [HRS16] for instance) is a particular case of Definition 2.2. The correctness of this assumption in several settings was studied in [FGK+19], such as strongly convex optimization. In Section 3, we will extend this condition to noisy algorithms.
In the context of topological generalization bounds, a different stability assumption was already used in [DDS23] to obtain generalization bounds with slightly simpler information-theoretic terms, but without removing the need for such terms. The main technical novelty of our study is to exploit stability to prove IT-free topological generalization bounds, as we present in the next section.
3 Main Theoretical Results
We now present our main theoretical contributions, which are expected topological generalization bounds on the generalization error $G_{S}(\mathcal{W}_{S})$ defined in Equation 3. Our core technical contribution is based on a variation of the hypothesis set stability of [FGK+19] (see Definition 2.2), which we present in Section 3.1. First, we specify some notation. Let $\mathrm{CL}(\mathbb{R}^{d})$ denote the closed subsets of $\mathbb{R}^{d}$ and $\Omega$ denote an arbitrary probability space, we see a learning algorithm as a mapping:
$\mathcal{A}:\bigcup_{n=1}^{\infty}\mathcal{Z}^{n}\times\Omega\longrightarrow\mathrm{CL}(\mathbb{R}^{d}),$ (7)
which we denote $\mathcal{A}(S,U)=\mathcal{W}_{S,U}$, where $S$ is the dataset and $U$ is an independent random variable representing the algorithm randomness. In practice, $\mathcal{W}_{S,U}$ is the trajectory of the optimizer in the parameter space, as specified in Section 1. This framework has been adopted by previous studies [HSKM22, DDS23]. Note that the dependence on $U$ will be omitted when the context is clear.
3.1 Random trajectory stability
The main drawback of Definition 2.2 is that it does not explicitly account for the algorithm randomness (e.g., batch indices), as opposed to the classical algorithmic stability [BE02, HRS16].
In order to address this issue, we propose a variant of the assumption of [FGK+19], accounting for the presence of randomness. The main difficulty that arises here is to give a meaning to the expressions such as $w\in\mathcal{W}_{S}$, when $\mathcal{W}_{S}$ is a random set. To do so, we define the notion of data-dependent selection [Mol17].
Definition 3.1.
A data-dependent selection of $\mathcal{W}_{S,U}$ is a deterministic mapping $\omega:\mathrm{CL}(\mathbb{R}^{d})\times\mathcal{Z}^{n}\to\mathbb{R}^{d}$ such that $\omega(\mathcal{W}_{S,U},S)\in\mathcal{W}_{S,U}$ almost surely with respect to the joint distribution of $(S,\mathcal{W}_{S,U})$.
We are mostly interested in the following “argmax selection”. In the rest of the paper, given $S_{1},S_{2}\in\mathcal{Z}^{n}$, we assume the existence of a random variable $\omega(\mathcal{W}_{S_{1}},S_{2})$ such that, almost surely:
$\omega(\mathcal{W}_{S_{1}},S_{2})\in\operatorname*{arg\,max}_{w\in\mathcal{W}_{S_{1}}}G_{S_{2}}(w).$ (8)
Note that the existence of such a random variable is ensured under mild conditions [Mol17]. Equipped with this definition, we now introduce our stability assumption.
Assumption 3.1 (Trajectory stability).
Algorithm $\mathcal{A}$ is $\beta_{n}$-trajectory-stable if for any $J\in\mathbb{N}^{\star}$ and any data-dependent selection $\omega$ of $\mathcal{W}_{S,U}$, there exists a mapping $\omega^{\prime}:\mathrm{CL}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that:
  • almost surely, for any $S$ and $w$, $\omega^{\prime}(W_{S},w)\in\mathcal{W}_{S}$
  • For all $z\in\mathcal{Z}$ and two datasets $S,S^{\prime}\in\mathcal{Z}^{n}$ differing by $J$ elements we have:
    $\mathbb{E}_{U}[|\ell(\omega(\mathcal{W}_{S,U},S),z)-\ell(\omega^{\prime}(\mathcal{W}_{S^{\prime},U},\omega(\mathcal{W}_{S,U},S)),z)|]\leq\beta_{n}J.$
In the absence of algorithmic randomness, that is, when each $U$ is constant, 3.1 is a particular case of Definition 2.2. Stability assumptions are usually enounced for neighboring datasets, here we present a variant with datasets differing by $J$ elements. These two variants are equivalent and we chose this presentation to simplify subsequent proofs and notations. The presence of the mapping $\omega^{\prime}$ might be intriguing: $\omega^{\prime}$ corresponds to the parameter denoted $w^{\prime}$ in Definition 2.2 and is denoted this way to make explicit its dependence on the different random variables.
The main advantage of 3.1 is to take the expectation over the algorithm randomness. A natural question arises as to whether this assumption can be satisfied in practice. We answer this question by the following lemma, which shows that 3.1 is implied by Definition 2.1.
Lemma 3.2.
[…]
This result shows that Definition 2.1 implies trajectory stability, in the sense of 3.1. While uniform argument stability differs from the classical notion of algorithmic stability, it can be noted that many stability-based studies rely on a Lipschitz assumption and prove uniform argument stability to derive algorithmic stability parameters [HRS16], as noted by [BFGT20]. Therefore, we see that the conditions of Lemma 3.2 can be satisfied in numerous cases, such as optimization of convex, smooth, and Lipschitz continuous functions [HRS16]. Under these conditions, it is obtained that the stability parameter is of order $\beta_{k}=\mathcal{O}(k/n)$, hence, yielding random trajectory stability with a parameter of order $\mathcal{O}(T^{2}/n)$, in the worst case. In the appendix, we present a corollary to Lemma 3.2 showing that projected SGD on a non-convex problem is $\mathcal{O}(1/n)$-trajectory-stable.
This finding is key to our approach. Indeed, combined with the topological generalization bounds as we do in the next section, a stability parameter of order $\mathcal{O}(1/n)$ will yield a complete bound of the form $\mathcal{O}(n^{-1/3}\sqrt{\mathbf{C}(\mathcal{W}_{S})})$, where $\mathbf{C}(\mathcal{W}_{S})$ is a topological complexity term.
3.2 Topological generalization bounds for trajectory-stable algorithms
We now leverage our random trajectory stability assumption to prove new topological generalization bounds without IT terms. We assume that the trajectories $\mathcal{W}_{S}$ are almost surely finite throughout, i.e., they correspond to the trajectory of the optimizer for a finite number of iterations.
As in previous work [DVDS24, ADS+24], our proofs are centered around the notion of Rademacher complexity [BM02] (formally defined in the appendix). One of our key technical contributions is deriving a Rademacher complexity bound for trajectory-stable algorithms. For the sake of simplicity, we postpone the technical details to the appendix and present our main results hereafter.
Our new topological generalization bounds rely on regularity conditions of the loss function (namely, Lipschitz continuity and boundedness), which we now present.
Assumption 3.3.
The loss function $\ell\colon\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}$ is $L$-Lipschitz continuous with respect to its first argument, i.e., for all $w,w^{\prime}\in\mathbb{R}^{d}$ and any $z\in\mathcal{Z}$, $|\ell(w,z)-\ell(w^{\prime},z)|\leq L\|w-w^{\prime}\|$.
Such Lipschitz assumptions were already used in previous topology-based studies [SSDE20, BLGS21, ALRS23]. Subsequent works [DDS23] mitigated these assumptions, albeit through the incorporation of intricate IT terms, which we avoid in our stability-based framework. As is typical in this literature [ADS+24, DVDS24, BLGS21], we also assume the boundedness of the loss function.
Assumption 3.4.
We assume that the loss $\ell$ is bounded in $[0,B]$, with $B>0$ being a constant.
We remark that this assumption holds; for instance, for the 00-$1$ loss. As demonstrated in [DDS23, ADS+24], the proposed theory remains experimentally valid even for unbounded losses. Our experimental results support extending this observation to our setting.
Equipped with these assumptions, we can now present our new stability-based topological generalization bounds. We present two results, depending on the topological complexity of interest. The first result is a topological bound based on $\alpha$-weighted lifetime sums.
Theorem 3.5.
Suppose that Assumptions 3.1, 3.3, and 3.4 hold and define $K_{n,\alpha}:=2(2L\sqrt{n}/B)^{\alpha}$. Without loss of generality [^5] Not making this assumption only yields changes in the absolute constants of the bound., assume that $\beta_{n}^{-2/3}$ is an integer divisor of $n$. For any $\alpha\in(0,1]$, we have:
$\mathbb{E}\bigg{[}\sup_{w\in\mathcal{W}_{S,U}}(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w))\bigg{]}\leq\beta_{n}^{1/3}\left(2+2B+2B\mathbb{E}\left[\sqrt{2\log(1+K_{n,\alpha}\mathbf{E}^{\alpha}_{\left\|\cdot\right\|}(\mathcal{W}_{S,U}))}\right]\right).$
The detailed proof is postponed to the appendix. By building on the results on positive magnitude of [ADS+24], this leads to the following positive magnitude-based generalization bound for trajectory-stable algorithms.
Theorem 3.6.
Suppose that Assumptions 3.1, 3.3, and 3.4 hold and define $s(\lambda):=\lambda L\beta_{n}^{-1/3}/B$ Without loss of generality, assume that $\beta_{n}^{-2/3}$ is an integer divisor of $n$. For any $\lambda>0$ we have:
$\mathbb{E}\bigg{[}\sup_{w\in\mathcal{W}_{S,U}}(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w))\bigg{]}\leq\beta_{n}^{1/3}\left(2+\lambda B+\frac{2B}{\lambda}\mathbb{E}\left[\log\mathbf{PMag}_{\left\|\cdot\right\|}\left(s(\lambda)\cdot\mathcal{W}_{S,U}\right)\right]\right).$
This result suggests that a theoretically justified choice of the magnitude scale would be $s(\lambda)\approx\beta_{n}^{-1/3}$, which amounts to $\Theta(n^{1/3})$ in the event that $\beta_{n}=\mathcal{O}(1/n)$. This is an interesting observation, as the results in [ADS+24] suggested an optimal scale parameter of order $s=\Theta(\sqrt{n})$.
Note moreover that it was also observed by that small values of $s(\lambda)$ can give satisfactory correlation with the generalization error, indicating that better choices of $\lambda$ might be available in certain cases. We leave the investigation of this matter to future studies.
In Theorems 3.5 and 3.6, we focus on two topological complexities: the weighted lifetime sums and the positive magnitude. This choice is motivated by the empirical results of [ADS+24], who demonstrated that they have better correlation with the generalization error compared to previously introduced dimension-based complexities [SSDE20]. Nevertheless, our approach can trivially recover bounds based on other complexities, such as the persistent homology dimension [BLGS21].
Our bounds show that for trajectory-stable algorithms, the topological complexities (i.e., lifetime sums and positive magnitude) have an important impact on the worst-case generalization error. As opposed to existing trajectory-dependent topological bounds [SSDE20, BLGS21, DDS23, ADS+24], our bounds do not feature any IT terms. In addition to replacing IT terms with a stability assumption, the stability parameter $\beta_{n}$ acts as a multiplying constant in the bound, hence increasing the impact of the topological complexity on the bound. Compared to the IT terms, $\beta_{n}$ is easily interpretable and can be expected to decrease with $n$ at a polynomial rate, as discussed above.
Despite being the first mutual information-free topological bounds, Theorems 3.5 and 3.6 have three main disadvantages: (i) we only provide expected bounds, (ii) we obtain a slightly worse convergence rate in $n$ and (iii) our framework only allows for Euclidean-based topological complexity, and we typically do not cover the case of the data-dependent pseudometric introduced by [DDS23]. Regarding (ii), our results should be compared with [DDS23, Theorem 3.8], which employed a stability-like assumption to obtain smaller IT terms than existing works at the time. They obtained a convergence rate that is similar to Theorems 3.5 and 3.6 but our results do not feature any IT term.
Stability parameters are expected to be affected by the sample size, hence, these results suggest to study the correlation between generalization error and topological complexity for several values of $n$, which forms a part of our empirical study.
4 Empirical Evaluations
We conduct all experiments with a Vision Transformer (ViT) [DBK+21] on CIFAR-100 [KH+09]. Additional implementation details are left for the appendix.
Experimental design
We follow the protocol of previous works [DDS23, ADS+24] and train models until convergence ($>93\%$ training accuracy), and use this checkpoint as a starting point for analysis of topological quantities. Our empirical analysis is then structured over a $4\times 4\times 5$ grid based on the following free variables: the number of training samples $n\in\{100,500,1000,5000,10000\}$, learning rates $\eta\in[10^{-6},10^{-4}]$, and batch sizes $b\in\{32,64,128,256\}$. We further investigated the stability parameter using a fixed learning rate of $10^{-6}$ and a batch size of $64$, testing across 10 different random seeds. These analyses are conducted without any further regularization techniques or learning rate scheduling to minimize the effect of confounding factors. All models are implemented in pytorch and trained with the ADAM optimizer.
Generalization error
As our theoretical results (see Section 3) provide bounds on the worst-case generalization gap $G_{S}(\mathcal{W}_{S,U})$, we seek to approximate this gap in practical settings. Prior work has typically used one of two proxies: the generalization gap at the final iteration, computed as final test risk -- final train risk [BLGS21], or the worst observed test risk over time, computed as worst-case test risk -- final train risk [ADS+24], where test risk was logged every $100$ iterations. To improve granularity, we assess both train and test risk every iteration and report worst-case (test risk - train risk) over all iterations.
Computing topological complexity
We employ the Krylov subspace approximation method to compute $\mathbf{PMag}(\mathcal{W}_{S,U})$, utilizing the preconditioned conjugate gradient (PCG) algorithm [Sal21]. Our computation of the $\mathbf{E}^{\alpha}(\mathcal{W}_{S,U})$ is based on the giotto-ph library [PHL+21], following the implementation of [ADS+24], using $\alpha=1$, which they argue is the most empirically relevant. In contrast to previous works [BLGS21, DDS23, ADS+24], which focus on detailed correlation analysis, our goal is to investigate how different topological quantities behave as a function of the sample size $n$, and how they relate instead of correlate to generalization gap. Hence, we refrain from introducing the granulated Kendall coefficients [JNM+19] and employ Pearson correlation coefficients to capture linear relationships. We present further correlation analysis in the appendix.
Distance matrix
The topological complexities that we consider in Section 3 are based on the Euclidean metric over the weight trajectory. Given a trajectory $\mathcal{W}_{S,U}$, the estimation of these quantities requires computing the distance matrix on all $\mathcal{W}_{S,U}$. To make this tractable, we uniformly sample $1500$ points from the full training trajectory, which consists of $5000$ iterations. Based on this sampled trajectory, we define the distance matrix using the Euclidean norm $\rho(w,w^{\prime}):=\left\|w-w^{\prime}\right\|$.
4.1 Results and Discussion
While previous works on topological generalization bounds [BLGS21, DDS23, ADS+24] analyzed the correlation between generalization and topological complexities, the intractability of IT components has thus far prevented a more complete analysis of the bounds. As our bound does not contain IT terms, relying primarily on the stability parameter defined by 3.1 and the topological quantities introduced in Theorems 3.5 and 3.6, all the terms they contain are amenable to numerical estimation, which is a unique feature among topological bounds. Next, we empirically characterize the interplay between stability and the topological complexities $\mathbf{E}^{\alpha}$ and $\mathbf{PMag}$.
Figure 1: Trajectory stability of the ViT for different training sample sizes $n$, for models with and without unseen injected samples $J$, for varying random seeds. The plot depicts models trained from a random initialization and converged pretrained models. We experiment with $\mathbf{Z}$ containing either validation or training points.
Stability Analysis
Our theoretical results depend on 3.1, which involves the stability of the trajectories $\mathcal{W}_{S,U}$ and provides a noisy variant of Definition 2.2. In our experiments, we assessed the assumption by a direct estimation of the stability coefficient. We focus on a condition similar to Definition 2.2 to avoid technical complications. Note that assuming the latter almost surely over $U$ is stronger than 3.1 so that the experiments can serve as an upper bound of Theorems 3.5 and 3.6.
To estimate the trajectory stability of each network, we replace a subset $\mathbf{J}$ of unseen samples for each training set $S\in\mathcal{Z}^{n}$ Section 4, where $|\mathbf{J}|=50$. After retraining the model with these perturbed datasets, we analyze the resulting trajectories $\mathcal{W}_{S,U}$ (original) and $\mathcal{W}_{S^{\prime},U}$ (perturbed). We quantify stability by computing the worst-case deviation in model losses across all iterations, analogous to a Haussdorf distance across the trajectories: $\max_{w\in\mathcal{W}_{S,U}}\min_{w^{\prime}\in\mathcal{W}_{S^{\prime},U}}\sup_{Z\in\mathbf{Z}}|\ell(w,Z)-\ell(w^{\prime},Z)|$, with $M:=\min\{n,500\}$ and $\mathbf{Z}:=\{Z_{1},\dots,Z_{M}\}$ a batch disjoint from $\mathbf{J}$, containing points either from the training or validation sets. Figure 1 shows the results for different sample sizes, where other hyperparameters are fixed. Note that we consider the loss difference over a batch $\mathbf{Z}$ of $m$ samples instead of taking the supremum over the whole dataset as in Definition 2.2. This procedures forms a tractable empirical substitute for Definition 2.2. To contextualize with 3.1 we report both the metric difference for train and validation sets.
The results are depicted in Figure 1. With increasing $n$, and thus, a decreasing proportion of the replaced samples, we observe that the per-iteration loss difference decreases, which is in accordance with established theoretical results [HRS16, BFGT20] and Lemma 3.2. Importantly, we observe a stable behavior when computing the loss differences for both the training and validation sets, suggesting that trajectory stability could be attained in practical settings with non-convex objectives.
This experiment shows that trajectory stability can be expected to hold in such settings. This observation might serve as a basis to understand the empirical success of previous studies [BLGS21, DDS23, ADS+24], in which the presence of the IT terms induced a discrepancy between theory and experiments. Our framework bridges this gap by rendering the bounds tractable. Moreover, Figure 1 suggests that stability holds when the experiment is started from a pre-trained weight, supporting previous findings that topological complexities correlate with $G_{S}(\mathcal{W}_{S,U})$ near convergence. ve that when the model is pretrained to full convergence, the per-iteration loss differences reflect a built-in stability, more so than for randomly initialized models.
Topological Complexity Analysis
Figure 2: Evolution of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ and $\mathbf{PMag}(100)$ with increasing sample size $n$, for learning rates $\eta=10^{-4}$ and $\eta=10^{-5}$.
Figure 3: Variation of $\mathbf{E}^{1}$ with sample size $n$. Pearson correlation coefficients $r$ are reported for each subgroup.
Our stability analysis shows that our assumptions hold both for large sample sizes $n$ as well as for randomly initialized and converged models. While the stability parameter is expected to decrease with increasing sample size, the behavior of the topological complexity as a function of $n$ has not been empirically studied in the literature. We analyze the behavior of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ and $\mathbf{PMag}(100\cdot\mathcal{W}_{S,U})$ in Figure 3. While our theory suggests letting the scale in $\mathbf{PMag}$ vary with $n$, we fix the scale to $100$, which is on the order of magnitude suggested by the theory. This allows us to isolate the effect that $n$ has on $\mathcal{W}_{S,U}$ when observing the variations of the reported quantity. We report these quantities for fixed learning rates to avoid confounding with large differences in generalization. We observe that, with increasing sample size $n$, the values of $\mathbf{E}^{1}$ and $\mathbf{PMag}$ increase and seem to stabilize for large values of $n$. This interesting observation leads to an interdependence between the sample size and the topology of the training trajectory, a notion that has not been reflected in prior works. We leave a detailed theoretical investigation as future work.
The behavior of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ as a function of $n$ might be puzzling at first, as we expect the generalization error and, hence, our bound to decrease with $n$. There is no contradiction here as the topological complexities $\mathbf{C}(\mathcal{W}_{S,U})$ should increase at most polynomially with $n$. Indeed, the positive magnitude is always bounded in terms of the number of iterations, and, based on persistent homology dimension and limiting arguments as $n\to\infty$, $\mathbf{E}^{\alpha}(\mathcal{W}_{S,U})$ also has sub-polynomial growth as $n\to\infty$ under mild assumptions (e.g., bounded trajectories), see for instance [BLGS21]. As argued in Section 3, the stability coefficient $\beta_{n}$ decreases polynomially in many cases. As Theorems 3.5 and 3.6 are logarithmic in $\mathbf{C}(\mathcal{W}_{S,U})$, this ensures that the bounds cannot be vacuous as $n\to\infty$.
In Figure 3, we observe that the sensitivity of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ with respect to the generalization gap (i.e., the slope of the affine regression of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ by the value of $G_{S}(\mathcal{W}_{S,U})$) increases when $n$ gets larger, which is what is predicted by our theory. Indeed, our main theorems assert that $\log\mathbf{E}^{1}(\mathcal{W}_{S,U})$ should be (approximately) of order at least $\beta_{n}^{-1/3}G_{S}(\mathcal{W}_{S,U})$, which amounts to $n^{1/3}G_{S}(\mathcal{W}_{S,U})$ in the event that $\beta_{n}=\Theta(1/n)$. Therefore, our experimental results strongly support Theorem 3.5.
Together, these observations unveil a strong coupling between stability and the topological complexity of the training trajectories. Moreover, the topological complexities increase for large $n$ and therefore become more relevant to the bound. The empirically observed coupling is supported by the theoretical results of Section 3, through the product of the stability parameter $\beta_{n}$ and the topological complexities.
5 Concluding Remarks
In this work, we present novel topological generalization bounds by making an assumption on the stability of training trajectories, demonstrating that this condition holds under practical settings. Based on this new framework, we obtain the first mutual information-free topological generalization bounds, which are, for the first time, tractable. This addresses one of the key theoretical issues in an emerging area of research. We empirically evaluate the interaction between the topological complexities, unveiling new connections between the sample size and the topology of the training trajectories. Our results show a strong coupling between stability and topological complexity and reveal compelling new behavior for varying sample size $n$. Empirical evidence supports our theoretical findings, suggesting a nuanced interplay between the stability-related and topological quantities that bound neural network generalization.
Limitations & future work
While our bounds are more tractable than previous approaches with IT terms, $\beta_{n}$ is still challenging to quantify in the general case. Exploring suitable strategies to approximate this term poses exciting opportunities. On the theoretical side, the topological complexities based on the data-dependent pseudometric used in [ADS+24] are still not compatible with our approach and incorporating them is an object of future studies.
Broader impact
Due to its generality, our framework of provable generalization bounds based on topological quantities and trajectory stability is not anticipated to have any specific negative impact.
Acknowledgements
The authors are grateful for the support of the Excellence Strategy of local and state governments, including computational resources of the LRZ AI service infrastructure provided by the Leibniz Supercomputing Center (LRZ), the German Federal Ministry of Education and Research (BMBF), and the Bavarian State Ministry of Science and the Arts (StMWK). T. B. was supported by a UKRI Future Leaders Fellowship (MR/Y018818/1) as well as a Royal Society Research Grant (RG/R1/241402). U.Ş. is partially supported by the French government under the management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR-19-P3IA-0001 (PRAIRIE 3IA Institute). B.D. and U.Ş. are partially supported by the European Research Council Starting Grant DYNASTY – 101039676.
Appendix A Trajectory Stability of Projected SGD on Nonconvex Problems
In Section 4.1, we empirically demonstrate that 3.1 is practical for modern neural networks like the Vision Transformer. What follows is a formal result demonstrating that 3.1 holds generally for projected Stochastic Gradient Descent (SGD) under a gradient Lipschitz condition.
Let us consider the projected SGD recursion on a loss function $\ell$, which can be non-convex. In particular, given a dataset $S=(z_{1},\dots,z_{n})\in\mathcal{Z}^{n}$, consider the following recursion:
$w_{k+1}=\Pi_{R}\left[w_{k}-\eta_{k}\nabla\ell(w_{k},z_{i_{k+1}})\right],$ (9)
where $i_{k}\in\{1,\dots,n\}$ is an index chosen at random at each iteration $k$ and $\Pi_{R}$ denotes the projection operator on the ball centered around 00 with radius $R$. Here, we fix the batch size to $1$ for clarity.
The next result establishes the trajectory-stability of SGD on Lipschitz and smooth (i.e., gradient Lipschitz) loss functions and is an adaptation of [HRS16, Theorem 3.12] to our setting.
Corollary A.1.
Fix $T\in\mathbb{N}$. Suppose that Assumption 3.3 holds, $\ell$ is gradient Lipschitz:
$\|\nabla\ell(w,z)-\nabla\ell(w^{\prime},z)\|\leq G\|w-w^{\prime}\|,$
for some $G<\infty$ and the step-size satisfies $\eta_{k}\leq c/k$ for some $c<1/G$. Then the SGD recursion on $K$ iterations in (9) is trajectory-stable with the stability parameter:
$\beta_{n}=\frac{4LR}{n-1}\left(\frac{L}{GR}\right)^{\frac{1}{Gc+1}}\sum_{k=1}^{T}k^{\frac{Gc}{Gc+1}}.$
Proof.
To show the trajectory-stability of SGD, we will first show that it is argument-stable in the sense of Definition 2.1, then we will use Lemma 3.2 to establish the trajectory stability.
To prove the argument stability, we almost exactly follow the steps in [HRS16, Theorem 3.12]. Let us consider a neighboring dataset $S^{\prime}=(z^{\prime}_{1},\dots,z^{\prime}_{n})\in\mathcal{Z}^{n}$ that differs from $S$ by one element, i.e., $z_{i}=z^{\prime}_{i}$ except one element. Consider the SGD recursion on $S^{\prime}$:
$w^{\prime}_{k+1}=\Pi_{R}\left[w^{\prime}_{k}-\eta_{k}\nabla\ell(w^{\prime}_{k},z^{\prime}_{i_{k+1}})\right].$
Since most of $S$ and $S^{\prime}$ coincide, at the first iterations it is highly likely that $w_{k}$ and $w^{\prime}_{k}$ will be identical, since they will pick the same data points. Let us denote the random variable $\kappa\geq 1$ that is the first time $z_{i_{\kappa}}\neq z^{\prime}_{i_{\kappa}}$. Consider $K\in\mathbb{N}$ with $1\leq K\leq T$. By using this construction, we have:
[…]
We now focus on the first term in the right-hand-side of the above equation:
[…]
$\leq$ $2R\>\mathbb{P}(\kappa\leq t_{0})$
$=$ $\frac{2Rt_{0}}{n}.$ (10)
The second parameter is upper-bounded in the proof of [HRS16, Theorem 3.12], which reads as follows:
[…]
By combining these two estimates, we obtain:
$\mathbb{E}_{U}[\|w_{K}-w^{\prime}_{K}\|]\leq$ $\frac{2Rt_{0}}{n}+\frac{2L}{G(n-1)}\left(\frac{K}{t_{0}}\right)^{Gc}.$
Choosing
$t_{0}=\left(\frac{L}{GR}\right)^{\frac{1}{Gc+1}}K^{\frac{Gc}{Gc+1}}$
yields
$\mathbb{E}_{U}[\|w_{K}-w^{\prime}_{K}\|]$ $\leq\frac{4R}{n-1}\left(\frac{L}{GR}\right)^{\frac{1}{Gc+1}}K^{\frac{Gc}{Gc+1}}$
$=:\delta_{K}.$
Since $\delta_{K}$ is independent of $S$ and $S^{\prime}$, this shows that the $k$-th iterate of SGD is $\delta_{k}$ argument-stable. Finally, we conclude by Lemma 3.2, that SGD is trajectory-stable with parameter
$\beta_{n}=L\sum_{k=1}^{K}\delta_{k}.$
This concludes the proof. ∎
Combining Corollary A.1 with our empirical evidence from Section 4.1, we demonstrate the practical validity of our assumption. In the subsequent analysis we provide the reasoning behind our main results, which are based on 3.1.
Appendix B Omitted Proofs
In this section, we present in full details the omitted proofs of our main results in Section 3.
Recent advances on topological generalization bounds are based on a data-dependent Rademacher complexity, leveraging PAC-Bayesian theory on random sets [DVDS24]. In our stability-based approach, the Rademacher complexity also plays a central role. We define the standard Rademacher complexity [SSBD14, BM02] as:
Definition B.1 (Rademacher complexity).
Let $\mathcal{W}$ be a set, $S\in\mathcal{Z}^{n}$, and $\boldsymbol{\epsilon}:=(\epsilon_{1},\dots,\epsilon_{n})\sim\mathcal{U}(\{-1,1\})^{\otimes n}$ be i.i.d. Rademacher random variables (i.e., centered Bernoulli variables). We define:
$\mathbf{Rad}_{S}(\mathcal{W}):=\mathbb{E}_{\boldsymbol{\epsilon}}\left[\sup_{w\in\mathcal{W}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\ell(w,z_{i})\right].$
Intuitively, the Rademacher complexity measures the ability of a model to fit random labels. Using this definition, we now establish our main result.
B.1 A key technical lemma: Rademacher bound for trajectory-stable algorithms
A key technical element of our approach is the following theorem, which presents an expected Rademacher complexity bound for the worst-case generalization error of trajectory-stable algorithms.
Theorem B.1.
Let $S\sim\mu_{z}^{\otimes n}$ be a dataset. Let $J,K\in\mathbb{N}^{\star}$ be such that $n=JK$, and $\tilde{S}_{J}\in\mathcal{Z}^{J}$ be independent of $S$. Assume that algorithm $\mathcal{A}$ satisfies 3.1 with parameter $\beta_{n}>0$, then:
$\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w))\right]\leq 2\mathbb{E}\left[\mathbf{Rad}_{\tilde{S}_{J}}(\mathcal{W}_{S,U})\right]+2J\beta_{n}.$
The proof is inspired by the data splitting technique used in [DDS23, Theorem 3.8], which we adapt to our purpose of using the trajectory stability assumption to remove the need for information-theoretic terms.
Proof.
In the following, we consider tree independent datasets $S,S^{\prime},S^{\prime\prime}\sim\mu_{z}^{\otimes n}$, with $S=(z_{1},\dots,z_{n})$ (and similar notation for $S^{\prime}$ and $S^{\prime\prime}$). Up to a slight change in the constant, we can without loss of generality consider $J,K\in\mathbb{N}^{\star}$ such that $n=JK$. We write the index et as a disjoint union as follows:
$\{1,\dots,n\}=\coprod_{k=1}^{K}N_{k},$
with for all $k$, $|N_{K}|=J$, where $\coprod$ denotes the disjoint union. We also introduce the notation $S_{k}\in\mathcal{Z}^{n}$, with:
$\forall i\in\{1,\dots,n\},~{}(S_{k})_{i}=\left\{\begin{array}[]{ll}z_{i}^{\prime},\quad i\in N_{k}\\ z_{i},\quad i\notin N_{k}\end{array}\right.$
Given two datasets $S_{1},S_{2}\in\mathcal{Z}^{n}$, we also consider a measurable mapping $w(S_{1},S_{2})$ (whose existence is satisfied in our setting [Mol17]) such that, almost surely:
$\omega(\mathcal{W}_{S_{1}},S_{2})\in\operatorname*{arg\,max}_{w\in\mathcal{W}_{S_{1}}}\left(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S_{2}}(w)\right).$
Recall that $\mathcal{W}_{S}\in\mathrm{CL}(\mathbb{R}^{d})$ (where $\mathrm{CL}(\mathbb{R}^{d})$ denotes the closed subsets of $\mathbb{R}^{d}$) is a data-dependent random trajectory, which we assume is a random compact set [Mol17]. Finally we recall the notation:
$G_{S}(\mathcal{W}_{S,U}):=\sup_{w\in\mathcal{W}_{S,U}}\left(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\right).$
We have:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]$ $=\mathbb{E}\left[\mathcal{R}(\omega(\mathcal{W}_{S,U},S))-\widehat{\mathcal{R}}_{S}(\omega(\mathcal{W}_{S,U},S)))\right]$
$=\frac{1}{n}\mathbb{E}\left[\sum_{k=1}^{K}\sum_{i\in N_{k}}\left(\ell(\omega(\mathcal{W}_{S,U},S),z_{i}^{\prime\prime})-\ell(\omega(\mathcal{W}_{S,U},S),z_{i})\right)\right].$
By Assumption 3.1 there exists a measurable function $\omega^{\prime}:\mathrm{CL}(\mathbb{R}^{d})\times\mathcal{Z}^{n}\to\mathbb{R}^{d}$ such that for any $z\in\mathcal{Z}$
$\mathbb{E}_{U}[|\ell(\omega(\mathcal{W}_{S,U},S),z)-\ell(\omega^{\prime}(\mathcal{W}_{S_{k},U},\omega(\mathcal{W}_{S,U}),S),z)|]\leq J\beta_{n}.$
We now apply the stability assumption on each term of the sum to get:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]$ $\leq 2J\beta_{n}+\frac{1}{n}\mathbb{E}\left[\sum_{k=1}^{K}\sum_{i\in N_{k}}\left(\ell(\omega^{\prime}(\mathcal{W}_{S_{k},U},\omega(\mathcal{W}_{S,U},S)),z_{i}^{\prime\prime})-\ell(\omega^{\prime}(\mathcal{W}_{S_{k},U},\omega(\mathcal{W}_{S,U},S)),z_{i})\right)\right]$
$\leq 2J\beta_{n}+\frac{1}{K}\mathbb{E}\left[\sum_{k=1}^{K}\sup_{w\in\mathcal{W}_{S_{k},U}}\frac{1}{J}\sum_{i\in N_{k}}\left(\ell(w,z_{i}^{\prime\prime})-\ell(w,z_{i})\right)\right]$
$\leq 2J\beta_{n}+\frac{1}{K}\mathbb{E}\left[\sum_{k=1}^{K}\sup_{w\in\mathcal{W}_{S,U}}\frac{1}{J}\sum_{i\in N_{k}}\left(\ell(w,z_{i}^{\prime\prime})-\ell(w,z_{i}^{\prime})\right)\right],$
where the last equation follows by linearity of the expectation independence by interchanging $z_{i}$ and $z_{i}^{\prime}$ for all $i\in N_{k}$. By noting that all the terms inside the first sum have the same distribution, we have:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]$ $\leq 2J\beta_{n}+\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\frac{1}{J}\sum_{i\in N_{k}}\left(\ell(w,z_{i}^{\prime\prime})-\ell(w,z_{i}^{\prime})\right)\right]$
$=2J\beta_{n}+\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\frac{1}{J}\sum_{i\in N_{1}}\left(\ell(w,z_{i}^{\prime\prime})-\ell(w,z_{i}^{\prime})\right)\right],$
As $\mathcal{W}_{S,U}$ is independent of $S^{\prime}$ and $S^{\prime\prime}$, we can apply the classical symmetrization argument. More precisely, let $(\epsilon_{1},\dots,\epsilon_{n})\sim\mathcal{U}(\{-1,1\})^{\otimes n}$ be i.i.d. Rademacher random variables (i.e., centered Bernoulli random variables), independent from the other random variable. We have:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]$ $\leq 2J\beta_{n}+\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\frac{1}{J}\sum_{i\in N_{1}}\epsilon_{i}\left(\ell(w,z_{i}^{\prime\prime})-\ell(w,z_{i}^{\prime})\right)\right]$
$\leq 2J\beta_{n}+\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\frac{1}{J}\sum_{i\in N_{1}}\epsilon_{i}\ell(w,z_{i}^{\prime\prime})\right]+\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\frac{1}{J}\sum_{i\in N_{1}}(-\epsilon_{i})\ell(w,z_{i}^{\prime})\right]$
$=2J\beta_{n}+2\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\frac{1}{J}\sum_{i\in N_{1}}\epsilon_{i}\ell(w,z_{i}^{\prime})\right].$
Therefore, by denoting $\tilde{S}_{J}=(z_{i}^{\prime})_{i\in N_{1}}$, we obtain by definition of the Rademacher complexity that:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]$ $\leq 2J\beta_{n}+2\mathbb{E}\left[\mathbf{Rad}_{\tilde{S}_{J}}(\mathcal{W}_{S,U})\right],$
which is the desired result. ∎
This result is based on leveraging our stability assumption to enable the use of the classical symmetrization argument [SSBD14]. Compared to previously proposed modified Rademacher complexity bounds, [FGK+19, SvEH+23], the above theorem has the advantage of directly involving the Rademacher complexity of the data-dependent hypothesis set $\mathcal{W}_{S}$, which tightens the bound and makes it significantly more empirically relevant.
Compared to previous studies [FGK+19, DVDS24], a drawback of our bound compared to these studies is that it only holds in expectation. Moreover, while we directly involve the data-dependent trajectory $\mathcal{W}_{S}$, our Rademacher $\mathbf{Rad}_{\tilde{S}}(\mathcal{W}_{S})$ term uses an independent copy of the dataset, hence, losing some data dependence compared to the term $\mathbf{Rad}_{S}(\mathcal{W}_{S})$ obtained by [DVDS24]. That being said, we show that our techniques still recover certain topological complexities, at the cost of requiring an additional Lipschitz continuity assumption. The Rademacher complexity term appearing in Theorem B.1 uses the data-dependent trajectory $\mathcal{W}_{S}$ and an independent dataset $\tilde{S}$ of size $J$. Following classical arguments [SSBD14, ADS+24], we expect this term to behave as $\mathcal{O}(1/\sqrt{J})$, hence, optimizing the value of $J$ yields a bound of order $\mathcal{O}(\beta_{n}^{1/3})$.
B.2 Proof of Theorems 3.5 and 3.6
Before presenting the proof of Theorems 3.5 and 3.6, we present two technical lemma, whose proof can be found in [ADS+24]. The first lemma relates the Rademacher complexity to the $\alpha$-weighted lifetime sums.
Lemma B.2.
Let $m\in\mathbb{N}^{\star}$, $S\in\mathcal{Z}^{m}$, and $\mathcal{W}\subset\mathbb{R}^{d}$ be a finite set. Assume that 3.4 and 3.3 hold. For any $\alpha\in(0,1]$, we have:
$\mathbf{Rad}_{S}(\mathcal{W})\leq\frac{B}{\sqrt{m}}+B\sqrt{\frac{2\log(1+K_{n,\alpha}\mathbf{E}^{\alpha}_{\left\|\cdot\right\|}(\mathcal{W}))}{m}},$
with $K_{n,\alpha}:=2(2L\sqrt{n}/B)^{\alpha}$.
Proof.
This result is a particular case of a bound by [ADS+24] as part of the proof of their Theorem 3.4. The assumptions of [ADS+24, Theorem 3.4] are satisfied by 3.4 and 3.3, and by considering the Euclidean distance of $\mathbb{R}^{d}$ as a pseudometric on $\mathcal{W}$. ∎
The second lemma relates the Rademacher complexity to the positive magnitude introduced by [ADS+24].
Lemma B.3.
Let $m\in\mathbb{N}^{\star}$, $S\in\mathcal{Z}^{m}$, and $\mathcal{W}\subset\mathbb{R}^{d}$ be a finite set. Assume that 3.4 and 3.3 hold. Then, we have for any $\lambda>0$ that:
$\mathbf{Rad}_{S}(\mathcal{W})\leq\frac{\lambda B^{2}}{2m}+\frac{1}{\lambda}\log\left(\mathbf{PMag}(L\lambda\mathcal{W})\right).$
This result was obtained in the proof of [ADS+24, Theorem 3.5] under a quite general Lipschitz regularity condition. For the sake of completeness, we present the proof in our particular case, as it might be more concrete for some readers.
Proof.
Let us fix $S:=(z_{1},\dots,z_{m})\in\mathcal{Z}^{m}$.
The proof follows the same strategy as the proof of [ADS+24, lemma B.13], with the difference that we use the definition of magnitude fo compact sets.
It was shown by [Mec15] that a finite set $\mathcal{W}$ has magnitude. More precisely, for any $\lambda>0$ there exists a vector $\beta_{s}:\mathcal{W}\to\mathbb{R}$, such that for any $a\in\mathcal{W}$ we have:
$1=\sum_{b\in\mathcal{W}}e^{-\lambda\left\|a-b\right\|}\beta_{\lambda}(b).$
Let $\epsilon=(\epsilon_{1},\dots,\epsilon_{m})\sim\mathcal{U}(\{-1,1\})^{\otimes m}$ be Rademacher random variables. By 3.3 we have for all $a\in\mathcal{W}$:
11 11 $\leq\sum_{b\in\mathcal{W}}e^{-\lambda L\left\|a-b\right\|}\beta_{\lambda L}(b)_{+}$
$\leq\sum_{b\in\mathcal{W}}\exp\left(-\frac{\lambda}{m}\sum_{i=1}^{m}|\ell(a,z_{i})-\ell(b,z_{i})|\right)\beta_{\lambda L}(b)_{+}$
$\leq\sum_{b\in\mathcal{W}}\exp\left(-\frac{\lambda}{m}\sum_{i=1}^{m}\epsilon_{i}(\ell(a,z_{i})-\ell(b,z_{i}))\right)\beta_{\lambda L}(b)_{+}.$
Therefore:
$\exp\left(\frac{\lambda}{m}\sum_{i=1}^{m}\epsilon_{i}\ell(a,z_{i})\right)\leq\sum_{b\in\mathcal{W}}\exp\left(\frac{\lambda}{m}\sum_{i=1}^{m}\epsilon_{i}\ell(b,z_{i})\right)\beta_{\lambda L}(b)_{+}.$ (11)
We apply it to the following choice of $a$:
$a=a(\epsilon):=\operatorname*{arg\,max}_{a\in\mathcal{W}}\frac{\lambda}{m}\sum_{i=1}^{m}\epsilon_{i}\ell(a,z_{i}).$
Thus, by Jensen’s inequality, we have:
$\mathbf{Rad}_{S}(\mathcal{W})$ $=\mathbb{E}_{\epsilon}\left[\max_{a\in\mathcal{W}}\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}\ell(a,z_{i})\right]$
$=\mathbb{E}_{\epsilon}\left[\frac{1}{m}\sum_{i=1}^{m}\epsilon_{i}\ell(a(\epsilon),z_{i})\right]$
$\leq\frac{1}{\lambda}\log\mathbb{E}_{\epsilon}\left[\exp\left(\frac{\lambda}{m}\sum_{i=1}^{m}\epsilon_{i}\ell(a(\epsilon),z_{i})\right)\right]$
$\leq\frac{1}{\lambda}\log\mathbb{E}_{\epsilon}\left[\sum_{b\in\mathcal{W}}\exp\left(\frac{\lambda}{m}\sum_{i=1}^{m}\epsilon_{i}\ell(b,z_{i})\right)\beta_{\lambda L}(b)_{+}\right].$
By 3.4, independence, and Hoeffding’s lemma, we obtain:
$\mathbf{Rad}_{S}(\mathcal{W})$ $\leq\frac{1}{\lambda}\log\sum_{b\in\mathcal{W}}\prod_{i=1}^{m}\exp\left(\frac{4B^{2}\lambda^{2}}{8m^{2}}\right)\beta_{\lambda L}(b)_{+}$
$=\frac{1}{\lambda}\log\left(\exp\left(\frac{B^{2}\lambda^{2}}{2m}\right)\sum_{b\in\mathcal{W}}\beta_{\lambda L}(b)_{+}\right)$
$=\frac{\lambda B^{2}}{2m}+\frac{1}{\lambda}\log\mathbf{PMag}(L\lambda\cdot\mathcal{W}),$
which is the desired result. ∎
We can finally prove Theorems 3.5 and 3.6.
See 3.5
Proof.
By Theorem B.1, we have (recall that $G_{S}$ has been defined in Equation 3):
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\mathbb{E}\left[\mathbf{Rad}_{\tilde{S}_{J}}(\mathcal{W}_{S,U})\right]+2J\beta_{n},$
with the notations of Theorem B.1 for $\tilde{S}_{J}$. Now we can apply Lemma B.2 to obtain that:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\left(\frac{B}{\sqrt{J}}+B\mathbb{E}\left[\sqrt{\frac{2\log(1+K_{n,\alpha}\mathbf{E}^{\alpha}_{\left\|\cdot\right\|}(\mathcal{W}_{S,U}))}{J}}\right]\right)+2J\beta_{n},$
We make the choice $J:=\beta_{n}^{-2/3}$, which we assumed for simplicity is an integer dividing $n$, so that it is a valid choice. This immediately gives the result. ∎
See 3.6
Proof.
By Theorem B.1, we have (recall that $G_{S}$ has been defined in Equation 3):
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\mathbb{E}\left[\mathbf{Rad}_{\tilde{S}_{J}}(\lambda L\mathcal{W}_{S,U})\right]+2J\beta_{n},$
with the notations of Theorem B.1 for $\tilde{S}_{J}$. Now we can apply Lemma B.2 to obtain that for any $\lambda>0$, we have:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\left(\frac{\lambda B^{2}}{2J}+\frac{1}{\lambda}\log\left(\mathbf{PMag}(\lambda L\mathcal{W}_{S,U})\right)\right)+2J\beta_{n}.$
We make the change of variable $\lambda\to\lambda\sqrt{J}/B$, which gives that for any $\lambda>0$, we have:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq\frac{\lambda B}{\sqrt{J}}+\frac{2B}{\lambda\sqrt{J}}\log\mathbf{PMag}(\frac{\lambda L\sqrt{J}}{B}\mathcal{W}_{S,U})+2J\beta_{n}.$
We make the choice $J:=\beta_{n}^{-2/3}$, which we assumed for simplicity is an integer dividing $n$, so that it is a valid choice. This immediately gives the result. ∎
B.3 Proof of Lemma 3.2
See 3.2
Proof.
Consider a data-dependent selection of $\mathcal{A}$, denoted $\omega(\mathcal{W},S)$, in the sense of Definition 3.1. We consider a measurable mapping $\omega^{\prime}:\mathrm{CL}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that for any closed set $\mathcal{W}\subset\mathbb{R}^{d}$ and any $w\in\mathbb{R}^{d}$, we have:
$\omega^{\prime}(\mathcal{W},w)\in\operatorname*{arg\,min}_{w^{\prime}\in\mathcal{W}}\left(\sup_{z\in\mathcal{Z}}|\ell(w,z)-\ell(w^{\prime},z)|\right).$
Such a mapping can be constructed in our setting where the sets are compact and the loss continuous. Consider two datasets $S,S^{\prime}$ in $\mathcal{Z}^{n}$ differing by $J$ elements. Let us denote:
$\varepsilon:=\sup_{z\in\mathcal{Z}}\mathbb{E}_{U}\left[\left|\ell(\omega(\mathcal{W}_{S,U},S),z)-\ell(\omega^{\prime}(\mathcal{W}_{S^{\prime},U},\omega(\mathcal{W}_{S,U},S)),z)\right|\right]$
We use the definition of $\omega^{\prime}$ and the Lipschitz-continuity of $\ell$ to obtain that:
$\varepsilon$ $\leq\sup_{z\in\mathcal{Z}}\mathbb{E}_{U}\left[\max_{1\leq k\leq K}\left|\ell(\mathcal{A}_{k}(S,U),z)-\ell(\omega^{\prime}(\mathcal{W}_{S^{\prime},U},\mathcal{A}_{k}(S,U)),z)\right|\right]$
$\leq\sup_{z\in\mathcal{Z}}\sum_{k=1}^{K}\mathbb{E}_{U}\left[\left|\ell(\mathcal{A}_{k}(S,U),z)-\ell(\omega^{\prime}(\mathcal{W}_{S^{\prime},U},\mathcal{A}_{k}(S,U)),z)\right|\right]$
$\leq\sum_{k=1}^{K}\mathbb{E}_{U}\left[\sup_{z\in\mathcal{Z}}\left|\ell(\mathcal{A}_{k}(S,U),z)-\ell(\omega^{\prime}(\mathcal{W}_{S^{\prime},U},\mathcal{A}_{k}(S,U)),z)\right|\right]$
$\leq\sum_{k=1}^{K}\mathbb{E}_{U}\left[\sup_{z\in\mathcal{Z}}\left|\ell(\mathcal{A}_{k}(S,U),z)-\ell(\mathcal{A}_{k}(S^{\prime},U),z)\right|\right]$
$\leq L\sum_{k=1}^{K}\mathbb{E}_{U}\left[\left\|\mathcal{A}_{k}(S,U)-\mathcal{A}_{k}(S^{\prime},U)\right\|\right]$
$=LJ\sum_{k=1}^{K}\delta_{k},$
where the last step follows by applying the stability assumption $J$ times and the triangle inequality. The result follows by definition of the random trajectory stability assumption. ∎
Appendix C Additional Experimental Details
This section expands on our experimental analyis, providing additional discussion of the methodology, implementation details, and technical setup.
C.1 Implementation Details
We implement training and validation loops in PyTorch Lightning Version $2.4$ using pytorch version $2.4.1$. Both models are trained using the AdamW optimizer with a cosine annealing learning rate scheduler.
Vision Transformer (ViT)
The ViT is configured for image classification on the CIFAR100 dataset. For the Vision Transformer on CIFAR100, we utilize a model with approximately 1.2 million parameters. This represents a reduction in size compared to the ViT experiments in [ADS+24], as the empirical analyses we conduct require substantially more experiments due to the expansion of training set size $n$ and thus storage of weight iterates.
The ViT processes 32x32 pixel images with 3 color channels, and a patch size of 4x4. The transformer backbone consists of 6 layers, 8 attention heads and an MLP dimension of 512. Dropout rates are set to 0.1 within the transformer blocks. Training is performed with a base learning rate of $0.0005$, decaying to a minimum of $10^{-5}$ via cosine annealing.
GraphSage
In addition to the main paper results, we analyze GraphSAGE [HYL17] on the graph classification task for MNISTSuperpixels [MBM+17]. It processes graphs where each node has a single input feature. These features are initially projected to a hidden dimension of 128. For graph-level predictions, node embeddings are aggregated using ‘mean’ global pooling, followed by a linear layer for classification into 10 classes. The model uses 2-dimensional positional encodings (representing x,y coordinates), which are added to the node features. We use 4 GraphSAGE convolutional layers with a dropout rate of 0.05, a base learning rate of 0.001, and weight decay of $0.0005$ to achieve an adequate generalization gap.
C.2 Stability Analysis
We first elaborate on our procedure to estimate trajectory stability in Section 4.1. The overall algorithm, a proxy to asses 3.1, is summarized in Algorithm 1. Stability experiments were performed with a fixed learning rate ($\eta$) and batch size ($b$), while varying random seeds, as described in Section 4.1. To ensure we capture two distinct training regimes, we analyze both randomly initialized and locally converged models; the latter is standard for the analysis of topological invariants [ADS+24]. We train each model for 500 additional iterations when analyzing trajectory stability.
The same setup is then conducted twice: once with the replacement set $J$ Section 4.1, and once without. Trajectory stability (3.1) is then estimated through the Hausdorf distance over each pair of trajectories where all model parameters are fixed except for the replacement set $J$.
Algorithm 1 Stability Parameter Estimation
1:Training set $S=\{z_{1},\dots,z_{n}\}$
2:Injection set $S_{J}=\{z^{\prime}_{1},\dots,z^{\prime}_{J}\}$ with $S_{J}\cap S=\emptyset$, and $J\leq n$
3:Estimate of stability parameter:
$\max_{w\in\mathcal{W}_{S,U}}\min_{w^{\prime}\in\mathcal{W}_{S^{\prime},U}}\sup_{Z\in\mathbf{Z}}|\ell(w,Z)-\ell(w^{\prime},Z)|$
4:procedure Training Phase
5:     Sample random indices $I=\{i_{1},\dots,i_{J}\}\subset\{1,\dots,n\}$
6:     Replace $\{z_{i}:i\in I\}$ with $S_{J}$ to form new dataset:
$S_{\text{new}}:=\left(S\setminus\{z_{i}:i\in I\}\right)\cup S_{J}$
7:     Train $T$ iterations on $S$ to obtain $\mathcal{W}_{S,U}=\{w_{1},\dots,w_{T}\}$
8:     Train $T$ iterations on $S_{\text{new}}$ to obtain $\mathcal{W}_{S_{\text{new}},U}=\{w^{\prime}_{1},\dots,w^{\prime}_{T}\}$
9:procedure Estimation Phase
10:     Set $M:=\min\{n,500\}$
11:     Sample $M$ indices $L=\{j_{1},\dots,j_{M}\}\subset\{1,\dots,n\}$
12:     Form evaluation set:
$\mathbf{Z}:=\{z_{j}\in S_{\text{new}}:j\in L\}\setminus S_{J}$
13:     Let $N:=|\mathbf{Z}|=M-J$
14:     Initialize matrices $A^{(1)},\dots,A^{(N)}\in\mathbb{R}^{T\times T}$
15:     for $i=1$ to $N$ do
16:         for $j=1$ to $T$ do
17:              for $l=1$ to $T$ do
18:                  $A^{(i)}_{j,l}:=|\ell(w_{j},Z_{i})-\ell(w^{\prime}_{l},Z_{i})|$
19:     Define $A\in\mathbb{R}^{T\times T}$ by:
$A_{j,l}:=\max_{i\in\{1,\dots,N\}}A^{(i)}_{j,l}$
20:     Compute vector $a\in\mathbb{R}^{T}$ where:
$a_{j}:=\min_{1\leq l\leq T}A_{j,l}$
21:     return Final estimate: $\max_{1\leq j\leq T}a_{j}$
C.3 Topological Complexity Analysis
In this section we detail our approach for weight trajectory analysis in terms of topological quantities, particularly experiments which demonstrate the relationship between stability and our topological complexity.
Following [ADS+24], we fine-tune each model from a local minimum for an additional 5000 iterations after training to convergence. We use a reduced parameter grid of $4\times 4$ for $\eta$ and $b$ ([ADS+24] use $6\times 6$), due to computational constraints arising from the varying the training set size $n$. Detailed hyperparameters are listed in Table 1.
Hyperparameter Vision Transformer GraphSage
Sample Size ($n$) $\{100,500,1000,5000,10000\}$$\{100,500,1000,5000,10000\}$
Learning Rate ($\eta$) $\{10^{-6},10^{-5},5\times 10^{-5},10^{-4}\}$$\{10^{-5},10^{-4},5\times 10^{-4},10^{-3}\}$
Batch Size ($b$) $\{32,64,128,256\}$$\{32,64,128,256\}$
Table 1: Summary of hyperparameter configurations for Vision Transformer and GraphSage models.
C.4 Hardware and Runtime
Models are trained on a cluster with 40GB A40/A100 GPUs; although the experiments are reproducible with lower memory requirements. For our grid experiments, computing topological features alongside hyperparameter variations (sample size, learning rate, and batch size) demanded approximately 57 GPU hours for GraphSage and 140 GPU hours for the Vision Transformer (ViT). In contrast, the stability analysis on the ViT required about 50 total GPU hours.
Appendix D Additional Empirical Results
In this section, we present additional empirical analyses and further substantiate the main paper analysis in Section 4.1 and Section 4.1.
D.1 Stability Estimation
In this section, we continue with the analysis of Section 4.1. We observe that, for both seen and unseen samples, an increase in sample size leads to a decrease in the estimated stability parameter (see Table 2). As shown in Section 4.1 and Table 2, the locally converged model demonstrates comparable stability between seen and unseen samples, with values similar to those of the randomly initialized model. This empirical evidence suggests that our trajectory stability assumptions are both theoretically sound and practically achievable in real-world applications.
Pretrain States Loss Difference Dataset Size ($n$) 100 500 1000 5000 10000 Randomly Initialized Seen 1.23 ± 0.22 0.56 ± 0.07 0.32 ± 0.04 0.07 ± 0.01 0.04 ± 0.01 Unseen 1.39 ± 0.16 0.60 ± 0.08 0.32 ± 0.04 0.07 ± 0.01 0.04 ± 0.00 Locally Converged Seen 1.20 ± 0.14 2.76 ± 0.38 2.37 ± 0.32 0.37 ± 0.07 0.36 ± 0.10 Unseen 4.60 ± 0.13 2.97 ± 0.23 1.59 ± 0.10 0.51 ± 0.10 0.28 ± 0.04
Table 2: Stability Analysis of Loss Differences Across Pretraining Iterations and Dataset Sizes for ViT Cifar100 (1.2M).
D.2 Topological Complexity Analysis
Building on the empirical analysis in Section 4.1, we elaborate further on how our topological complexities evolve with dataset size $n$.
Vision Transformer
Table 3 reports topological quantities for multiple scaling parameters of the positive magnitude, as well as $\mathbf{E}^{1}$. We observe a consistent trend across all topological features, mirroring the patterns identified in Section 4.1. This suggests that the observed behavior is not isolated to a specific topological quantity.
GraphSage
In addition to the experiments on the Vision Transformer, we validate our observation on a conceptually different model. We conduct the same experiment for GraphSage, as introduced in Section C.1. In these experiments, we similarly observe an increase in the magnitude of topological complexities with increasing sample size $n$. Figure 5 depicts a representative example.
While a similar behavior of increasing slope is apparent for GraphSage, the overall pattern is less apparent. This can be partially attributed to the larger discrepancies in the generalization gap, which are more difficult to modulate precisely by varying $\eta$ and $b$.
Nevertheless, as seen in the relationship between the generalization gap and $\mathbf{E}^{1}$ Figure 5, the slope similarly trends upwards with sample size $n$.
We additionally report positive magnitude for scaling that is theoretically supported ($3]{n}$), and the fixed scale $\mathbf{PMag}(100\cdot\mathcal{W_{S}})$. For the theoretically supported scales of $\mathbf{PMag}(\sqrt[3]{n}\cdot\mathcal{W_{S}})$, we observe a similar tendency of an increase in slope. Because the positive magnitude intuitively represents the effective number of distinct points in a space at a given scale $\lambda$, excessively large values of $\lambda$ can produce distorted and uninformative quantities. This occurs as fine-grained structural details are lost when the scale exceeds meaningful topological variations.
Figure 4: Topological invariants $\mathbf{E}^{1}$ and PMag for GraphSage on MNISTSuperpixel. PMag is depicted for the fixed scale of 100. The mean and std. err. over the batch sizes are shown for learning rates $10^{-4},10^{-3}$.
Figure 5: Behavior of $\mathbf{E}^{1}$ over $n$ for GraphSage on MNISTSuperpixel. Pearson correlation is reported per training subset $n$.
ViT Cifar100 Measure Dataset Size ($n$) 100 500 1000 5000 10000 $\mathbf{PMag}(\sqrt[3]{100}\cdot\mathcal{W_{S}})$ $\tau$ 0.733 0.767 0.667 0.717 0.633 $r$ 0.971 0.914 0.912 0.928 0.822 Slope 47.225 97.009 172.533 384.941 511.956 $\mathbf{PMag}(\sqrt[3]{500}\cdot\mathcal{W_{S}})$ $\tau$ 0.717 0.800 0.700 0.767 0.633 $r$ 0.973 0.921 0.924 0.940 0.834 Slope 87.332 207.714 371.531 790.149 1043.039 $\mathbf{PMag}(\sqrt[3]{1000}\cdot\mathcal{W_{S}})$ $\tau$ 0.717 0.817 0.700 0.750 0.633 $r$ 0.974 0.930 0.931 0.944 0.838 Slope 113.900 278.563 493.025 1021.470 1342.187 $\mathbf{PMag}(\sqrt[3]{5000}\cdot\mathcal{W_{S}})$ $\tau$ 0.700 0.800 0.733 0.767 0.650 $r$ 0.975 0.952 0.944 0.948 0.843 Slope 205.402 499.767 851.550 1655.984 2151.617 $\mathbf{PMag}(\sqrt[3]{10000}\cdot\mathcal{W_{S}})$ $\tau$ 0.700 0.783 0.733 0.750 0.633 $r$ 0.975 0.960 0.947 0.947 0.842 Slope 259.753 616.699 1030.246 1948.297 2519.328 $\mathbf{E}^{\alpha}$ $\tau$ 0.750 0.750 0.650 0.683 0.533 $r$ 0.976 0.957 0.942 0.948 0.842 Slope 33.652 73.124 123.392 247.425 325.163 $\mathbf{PMag}(100\cdot\mathcal{W_{S}})$ $\tau$ 0.783 0.767 0.650 0.700 0.650 $r$ 0.947 0.981 0.921 0.893 0.794 Slope 814.032 1400.576 2013.602 3177.603 4022.352
Table 3: We report the Kendall correlation coefficient $\tau$, the Pearson correlation coefficient $r$, and the slope for the Vision Transformer model on Cifar100. Bolded values denote the theoretically optimal scaling parameter of $\mathbf{PMag}$ corresponding to sample size n.
GraphSage Measure Dataset Size ($n$) 100 500 1000 5000 10000 $\mathbf{PMag}(\sqrt[3]{100}\cdot\mathcal{W_{S}})$ $\tau$ 0.850 0.533 0.683 0.667 0.500 $r$ 0.896 0.682 0.924 0.566 0.233 Slope 39.653 62.366 48.664 124.134 75.402 $\mathbf{PMag}(\sqrt[3]{500}\cdot\mathcal{W_{S}})$ $\tau$ 0.867 0.617 0.750 0.667 0.517 $r$ 0.897 0.691 0.929 0.564 0.241 Slope 67.973 107.236 79.759 226.185 142.980 $\mathbf{PMag}(\sqrt[3]{1000}\cdot\mathcal{W_{S}})$ $\tau$ 0.900 0.617 0.767 0.717 0.533 $r$ 0.898 0.692 0.928 0.570 0.250 Slope 85.249 134.828 96.132 276.827 180.090 $\mathbf{PMag}(\sqrt[3]{5000}\cdot\mathcal{W_{S}})$ $\tau$ 0.883 0.583 0.800 0.683 0.517 $r$ 0.904 0.691 0.914 0.598 0.280 Slope 142.404 225.500 140.712 403.594 283.260 $\mathbf{PMag}(\sqrt[3]{10000}\cdot\mathcal{W_{S}})$ $\tau$ 0.883 0.583 0.800 0.683 0.467 $r$ 0.908 0.690 0.900 0.617 0.298 Slope 176.644 279.671 163.426 461.667 335.112 $\mathbf{E}^{\alpha}$ $\tau$ 0.883 0.550 0.700 0.483 0.367 $r$ 0.921 0.683 0.922 0.571 0.279 Slope 19.361 30.711 26.662 78.674 56.366 $\mathbf{PMag}(100\cdot\mathcal{W_{S}})$ $\tau$ 0.917 0.567 0.767 0.633 0.483 $r$ 0.942 0.696 0.729 0.796 0.541 Slope 635.358 1092.277 433.767 941.365 855.726
Table 4: We report the Kendall correlation coefficient $\tau$, the Pearson correlation coefficient $r$, and the slope for the GraphSage model on MNISTSuperpixel. Bolded values denote the theoretically optimal scaling parameter of $\mathbf{PMag}$ corresponding to sample size n.